# On the Value of Cross-Modal Misalignment in Multimodal Representation Learning

### v4

@@ C.2 Additional Identification Results
As shown in Figure 9, the results of predicting true latent semantic variables under various perturbation bias settings exhibit similar trends to those observed under selection bias. Modality-specific variables are consistently discarded, unbiased semantic variables are faithfully block-identified, and misaligned semantic variables become partially identifiable in scenarios with dependent latent variables—demonstrating a consistent pattern across both image and text modalities. These findings further support and reinforce our theoretical analysis.
@@
We further analyze the linearity of the identified semantic representations by reporting the $R^{2}$ scores obtained from linear regression applied to the learned features for predicting the true latent variables. As shown in Figure 10 (selection bias) and Figure 11 (perturbation bias), the performance of linear regression closely mirrors that of nonlinear regression reported in Figure 3. This strong correspondence suggests that the relationship between the learned representations and the true latent semantic variables is approximately linear, indicating that the identified representation subspace is nearly linear.
@@
First, we investigate the impact of assigning an incorrect representation dimension. Specifically, we consider a scenario in which the true dimension of the selected semantic variables is $3$ (with selection $\mathbb{I}_{\theta}=[3]$), but we intentionally set the representation dimension to $5$. As shown in Table 4, in the independent case, all selected semantic variables are successfully preserved, while omitted semantic variables are effectively discarded. In contrast, the dependent scenario yields significantly different patterns of $R^{2}$ scores compared to those obtained using the correct representation dimension (see Figures 10 and 3). These results suggest that redundant representation dimensions tend to encode exogenous noise, potentially introducing unnecessary complexity into the learned representations.
Second, we explore the joint effect of selection and perturbation biases by defining a scenario with selection bias $\mathbb{I}_{\theta}=[8]$ and perturbation bias $\mathbb{I}_{\rho}=[2]$. Results presented in Table 5 demonstrate that when both biases coexist, their effects on semantic identification remain consistent: semantic variables that are either omitted or perturbed are discarded, while unbiased semantic variables—those that are selected and yet unperturbed—are reliably preserved in the learned representations.
Together with previous results, these findings further reinforce our theoretical conclusions in Thm. 4.1.
@@
C.3 Additional Downstream Results
We further report downstream task performance under varying perturbation bias settings. Specifically, the preserved semantic variables are sequentially reversed—starting from semantic index $10$ and incrementally expanding until the full semantic set is included. The results shown in Figure 12 indicate that, in general, semantic variables critical to downstream tasks must be preserved in the learned representations to achieve high performance. This observation holds across both independent and dependent latent semantic settings.
@@ C.3 Additional Downstream Results
Appendix D Experiment Details on MPI3D-Complex Dataset
We provide additional details on the MPI3D-Complex dataset that are not fully covered in § 5.2. In § D.1, we comprehensively describe the experimental setup, including a dataset overview, the selection and perturbation bias configurations used to generate text, model architecture, and training parameters. In § D.2, we present additional results, including the use of a linear classifier for predicting latent factors and an ablation study on encoder dimensionality.
D.1 Detailed Experimental Setup
@@
MPI3D-Complex [18] contains 460,800 real-world images of resolution $64\times 64\times 3$, spanning all combinations of seven mutually independent, discrete factors (see Table 6). We designate horizontal and vertical positions (hori., vert.) as image-specific due to their low-level visual nature, while treating the remaining five as semantic variables for representation learning and evaluation purposes.
@@
Text descriptions are generated from the ground-truth semantic attributes using content-word mappings (see Table 6). Under unbiased settings, all five semantic attributes are included. For text-specific variation, a generation template is chosen via a discrete latent variable $m_{t}\sim\mathrm{Uniform}([3])$, each corresponding to a different human-written sentence structure.
@@
Object shape (shape) $\mathrm{Uniform}(\{0,\dots,3\})$ coffee-cup, tennis-ball, croissant, beer-cup
@@
We generate text for each image under various selection and perturbation bias settings to investigate the effects of misalignment. To ensure computational tractability—since exhaustively enumerating all possible configurations is both infeasible and unnecessary—we adopt a progressively incremental strategy for introducing biases into the latent semantic variables, as outlined in Table 7.
In selection bias settings (where the perturbable set $\mathbb{I}_{\rho}$ is empty), textual descriptions include only the content words corresponding to a subset of the true image semantic variables. We define five incremental settings: 1: {color}, 2: {color, shape}, 3: {color, shape, size}, 4: {color, shape, size, cam.}, 5: all five attributes. The specific text generation templates associated with $\mathbf{g}_{t^{\scriptscriptstyle(\theta)}}$ under each selection setting are detailed in Table 8.
[…]
@@
For each setting, the dataset is partitioned into training, evaluation, and test subsets in a fixed ratio of $44{,}720:23{,}040:23{,}040$. Across all configurations, we train the image and text encoders for $200{,}000$ steps using the training subset, averaging results over three random seeds. The training objective is the multimodal contrastive loss $\mathcal{L}_{\text{\tiny MMCL}}$, defined in Eq. 1, and optimized using the Adam optimizer with an initial learning rate of $1\times 10^{-5}$, a batch size of 256, and a temperature parameter $\tau=1.0$.
@@
After training the representations, we freeze the encoders and train both a linear classifier (logistic regression) and a nonlinear classifier (a two-layer MLP with ReLU activation) for each setting and each image latent factor. Classifiers are trained on the evaluation subset for $10{,}000$ steps.
@@ D.2 Additional Results
We evaluate the linear separability of the learned representations by reporting MCC scores for predicting each latent factor using a linear classifier. As shown in Figure 13, and in comparison to the nonlinear results in Figure 6, the findings reveal that certain latent semantic variables—most notably size—are not linearly embedded in the learned representation space when training image-text pairs with MMCL.
@@
As shown in Figure 14, in contrast to training with purely numerical data, learning from image-text data exhibits sensitivity to the choice of encoding size. Specifically, when the encoding size is set to 1—exactly matching the number of perfectly aligned semantic dimensions—the image encoder tends to be under-optimized, resulting in high variance across runs. However, increasing the encoding size leads to more stable and reliable performance.
@@
Appendix E Experiment Details on Causal3DIdent Dataset
We provide additional details on the Causal3DIdent dataset that are not fully covered in § 5.3. In § E.1, we offer a comprehensive description of the experimental setup, including the image and text latent factors, the image generation process, the design of selection and perturbation bias settings for text generation, and training configurations. In § E.2, we present supplementary results, including analyses of the learned text representations and assessments of the linearity of the learned representations.
E.1 Detailed Experimental Setup
@@
These include 3 discrete variables—object shape (shape), and object positions along the horizontal (x_pos) and vertical (y_pos) axes—and 7 continuous variables: object color (color), spotlight position (s_pos) and color (s_color), background color (b_color), and the three object rotation angles (alpha, beta, gamma). We treat the rotation angles (alpha, beta, gamma) as image-specific latent variables, while the remaining factors are considered semantic latent variables, structured according to the causal graph shown in Figure 15.
We synthesize $80{,}000$ samples for MMCL training, $10{,}000$ samples for classifier or regressor training, and another $10{,}000$ samples for test-time evaluation. Images are rendered at a resolution of $128\times 128\times 3$.
@@
To introduce text-specific variation, we employ five manually designed templates to generate text from the latent factors under each bias setting, adapting the text rendering pipeline from [12]. A complete list of latent factors and their types is provided in Table 9.
@@
shape $\mathrm{Uniform}(\{0,\dots,6\})$ $\mathrm{Uniform}(\{0,\dots,6\})$ teapot, hare, dragon, cow, armadillo, horse, head
@@
s_pos $\mathrm{Uniform}([0,1])$ $\mathrm{Uniform}([0,1])$ northwest, northeast, center, southwest, southeast
@@
Following the MPI3D-Complex experiments, we explore a series of incrementally increasing selection and perturbation bias configurations to introduce varying degrees and types of cross-modal misalignment. These settings enable a systematic investigation of how different forms of alignment impact representation learning. Each configuration is indexed using circled numerals and summarized in Table 10.
@@
Table 11 provides the text generation templates associated with each selection setting. Representative image–text pairs generated under different selection and perturbation bias configurations are shown in Figure 16.
@@
Across all experimental settings, we train the image and text encoders for $100{,}000$ steps on the training subset, using three different random seeds to ensure robustness. The training objective is the multimodal contrastive loss $\mathcal{L}_{\text{\tiny MMCL}}$, defined in Eq. 1, and optimized using the Adam optimizer with an initial learning rate of $1\times 10^{-5}$, a batch size of 256, and a temperature parameter $\tau=1.0$.
@@
After training the representations, we freeze the encoders and, for each bias setting, train both a linear classifier (logistic regression) and a nonlinear classifier (a two-layer MLP with ReLU activation) for each discrete latent factor. Similarly, for continuous latent factors, we train both a linear regressor and a nonlinear regressor (a two-layer MLP with ReLU activation). All classifiers and regressors are trained for $10{,}000$ steps using the evaluation subset.
@@ E.2 Additional Results
We now turn to the analysis of the text representations learned by MMCL. As shown in Figure 17, we observe patterns similar to those found in the image modality. In particular, discrete latent semantic variables—such as shape, x_pos, and y_pos—are reliably identified, provided they are unbiased and consistently aligned across modalities. A notable case is s_pos, which is a continuous latent factor in both modalities but is mapped to only five discrete tokens in the text observations, rendering the text generation process non-invertible. Despite this lossy transformation, the model achieves a relatively high $R^{2}$, suggesting that the learned text representations remain strongly influenced by the alignment objective, even when semantic information is partially lost.
@@
We assess the linear separability of the learned image and text representations by evaluating the performance of linear classifiers and regressors trained to predict each latent factor. As shown in Figure 18 and Figure 19, and in comparison to the nonlinear prediction results, the findings indicate that not all latent semantic variables are linearly embedded in the representations learned through MMCL. In particular, the semantic factor color consistently demonstrates poor linear predictability, suggesting that it is encoded nonlinearly in both modalities. Conversely, factors such as x_pos and y_pos exhibit strong linear separability in certain settings, indicating that these semantic variables are more directly captured in the latent space of both the image and text encoders. Overall, these results suggest that the linearity of the learned representations is factor-dependent and shaped by both modality-specific encoding strategies and the underlying structure of the input data—highlighting an important direction for future investigation.
Appendix F Computation Resources
@@
Appendix G Discussions

### v5

@@ D.2 Additional Identification Results
As shown in Figure˜11, the results of predicting true latent semantic variables under various perturbation bias settings exhibit similar trends to those observed under selection bias. Modality-specific variables are consistently discarded, unbiased semantic variables are faithfully block-identified, and misaligned semantic variables become partially identifiable in scenarios with dependent latent variables—demonstrating a consistent pattern across both image and text modalities. These findings further support and reinforce our theoretical analysis.
@@
We further analyze the linearity of the identified semantic representations by reporting the $R^{2}$ scores obtained from linear regression applied to the learned features for predicting the true latent variables. As shown in Figure˜12 (selection bias) and Figure˜13 (perturbation bias), the performance of linear regression closely mirrors that of nonlinear regression reported in Figure˜3. This strong correspondence suggests that the relationship between the learned representations and the true latent semantic variables is approximately linear, indicating that the identified representation subspace is nearly linear.
@@
First, we investigate the impact of assigning an incorrect representation dimension. Specifically, we consider a scenario in which the true dimension of the selected semantic variables is $3$ (with selection $\mathbb{I}_{\theta}=[3]$), but we intentionally set the representation dimension to $5$. As shown in Table˜4, in the independent case, all selected semantic variables are successfully preserved, while omitted semantic variables are effectively discarded. In contrast, the dependent scenario yields significantly different patterns of $R^{2}$ scores compared to those obtained using the correct representation dimension (see Figures˜12 and 3). These results suggest that redundant representation dimensions tend to encode exogenous noise, potentially introducing unnecessary complexity into the learned representations.
Second, we explore the joint effect of selection and perturbation biases by defining a scenario with selection bias $\mathbb{I}_{\theta}=[8]$ and perturbation bias $\mathbb{I}_{\rho}=[2]$. Results presented in Table˜5 demonstrate that when both biases coexist, their effects on semantic identification remain consistent: semantic variables that are either omitted or perturbed are discarded, while unbiased semantic variables—those that are selected and yet unperturbed—are reliably preserved in the learned representations.
Together with previous results, these findings further reinforce our theoretical conclusions in Thm.˜4.1.
@@
D.3 Additional Downstream Results
We further report downstream task performance under varying perturbation bias settings. Specifically, the preserved semantic variables are sequentially reversed—starting from semantic index $10$ and incrementally expanding until the full semantic set is included. The results shown in Figure˜14 indicate that, in general, semantic variables critical to downstream tasks must be preserved in the learned representations to achieve high performance. This observation holds across both independent and dependent latent semantic settings.
@@ D.3 Additional Downstream Results
Appendix E Experiment Details on MPI3D-Complex Dataset
We provide additional details on the MPI3D-Complex dataset that are not fully covered in §˜5.2. In §˜E.1, we comprehensively describe the experimental setup, including a dataset overview, the selection and perturbation bias configurations used to generate text, model architecture, and training parameters. In §˜E.2, we present additional results, including the use of a linear classifier for predicting latent factors and an ablation study on encoder dimensionality.
E.1 Detailed Experimental Setup
@@
MPI3D-Complex [19] contains 460,800 real-world images of resolution $64\times 64\times 3$, spanning all combinations of seven mutually independent, discrete factors (see Table˜6). We designate horizontal and vertical positions (hori., vert.) as image-specific due to their low-level visual nature, while treating the remaining five as semantic variables for representation learning and evaluation purposes.
@@
Text descriptions are generated from the ground-truth semantic attributes using content-word mappings (see Table˜6). Under unbiased settings, all five semantic attributes are included. For text-specific variation, a generation template is chosen via a discrete latent variable $m_{t}\sim\mathrm{Uniform}([3])$, each corresponding to a different human-written sentence structure.
@@
Object shape (shape) $\mathrm{Uniform}(\{0,\dots,3\})$
coffee-cup, tennis-ball, croissant,
beer-cup
@@
We generate text for each image under various selection and perturbation bias settings to investigate the effects of misalignment. To ensure computational tractability—since exhaustively enumerating all possible configurations is both infeasible and unnecessary—we adopt a progressively incremental strategy for introducing biases into the latent semantic variables, as outlined in Table˜7.
In selection bias settings (where the perturbable set $\mathbb{I}_{\rho}$ is empty), textual descriptions include only the content words corresponding to a subset of the true image semantic variables. We define five incremental settings: 1: {color}, 2: {color, shape}, 3: {color, shape, size}, 4: {color, shape, size, cam.}, 5: all five attributes. The specific text generation templates associated with $\mathbf{g}_{t^{\scriptscriptstyle(\theta)}}$ under each selection setting are detailed in Table˜8.
In perturbation bias settings (where all semantic indices are selected), we apply random substitutions to a subset $\mathbb{I}_{\rho}$. For each factor in $\mathbb{I}_{\rho}$, its text value is replaced with a randomly sampled alternative with probability 0.9. Five perturbation configuration are used: 1: $\emptyset$, 2: {back.}, 3: {cam., back.}, 4: {size, cam., back.}, 5: {shape, size, cam., back.}.
@@
For each setting, the dataset is partitioned into training, evaluation, and test subsets in a fixed ratio of 44,720:23,040:23,04044{,}720:23{,}040:23{,}040. Across all configurations, we train the image and text encoders for 200,000200{,}000 steps using the training subset, averaging results over three random seeds. The training objective is the multimodal contrastive loss $\mathcal{L}_{\text{\tiny MMCL}}$, defined in Eq.˜1, and optimized using the Adam optimizer with an initial learning rate of $1\times 10^{-5}$, a batch size of 256, and a temperature parameter $\tau=1.0$.
@@
After training the representations, we freeze the encoders and train both a linear classifier (logistic regression) and a nonlinear classifier (a two-layer MLP with ReLU activation) for each setting and each image latent factor. Classifiers are trained on the evaluation subset for 10,00010{,}000 steps.
@@ E.2 Additional Results
We evaluate the linear separability of the learned representations by reporting MCC scores for predicting each latent factor using a linear classifier. As shown in Figure˜15, and in comparison to the nonlinear results in Figure˜6, the findings reveal that certain latent semantic variables—most notably size—are not linearly embedded in the learned representation space when training image-text pairs with MMCL.
@@
As shown in Figure˜16, in contrast to training with purely numerical data, learning from image-text data exhibits sensitivity to the choice of encoding size. Specifically, when the encoding size is set to 1—exactly matching the number of perfectly aligned semantic dimensions—the image encoder tends to be under-optimized, resulting in high variance across runs. However, increasing the encoding size leads to more stable and reliable performance.
@@
Appendix F Experiment Details on Causal3DIdent Dataset
We provide additional details on the Causal3DIdent dataset that are not fully covered in §˜5.3. In §˜F.1, we offer a comprehensive description of the experimental setup, including the image and text latent factors, the image generation process, the design of selection and perturbation bias settings for text generation, and training configurations. In §˜F.2, we present supplementary results, including analyses of the learned text representations and assessments of the linearity of the learned representations.
F.1 Detailed Experimental Setup
@@
These include 3 discrete variables—object shape (shape), and object positions along the horizontal (x_pos) and vertical (y_pos) axes—and 7 continuous variables: object color (color), spotlight position (s_pos) and color (s_color), background color (b_color), and the three object rotation angles (alpha, beta, gamma). We treat the rotation angles (alpha, beta, gamma) as image-specific latent variables, while the remaining factors are considered semantic latent variables, structured according to the causal graph shown in Figure˜17.
We synthesize 80,00080{,}000 samples for MMCL training, 10,00010{,}000 samples for classifier or regressor training, and another 10,00010{,}000 samples for test-time evaluation. Images are rendered at a resolution of $128\times 128\times 3$.
@@
To introduce text-specific variation, we employ five manually designed templates to generate text from the latent factors under each bias setting, adapting the text rendering pipeline from [13]. A complete list of latent factors and their types is provided in Table˜9.
@@
shape $\mathrm{Uniform}(\{0,\dots,6\})$ $\mathrm{Uniform}(\{0,\dots,6\})$
teapot, hare, dragon, cow,
armadillo, horse, head
@@
s_pos $\mathrm{Uniform}([0,1])$ $\mathrm{Uniform}([0,1])$
northwest, northeast, center,
southwest, southeast
@@
Following the MPI3D-Complex experiments, we explore a series of incrementally increasing selection and perturbation bias configurations to introduce varying degrees and types of cross-modal misalignment. These settings enable a systematic investigation of how different forms of alignment impact representation learning. Each configuration is indexed using circled numerals and summarized in Table˜10.
@@
Table˜11 provides the text generation templates associated with each selection setting. Representative image–text pairs generated under different selection and perturbation bias configurations are shown in Figure˜18.
@@
Across all experimental settings, we train the image and text encoders for 100,000100{,}000 steps on the training subset, using three different random seeds to ensure robustness. The training objective is the multimodal contrastive loss $\mathcal{L}_{\text{\tiny MMCL}}$, defined in Eq.˜1, and optimized using the Adam optimizer with an initial learning rate of $1\times 10^{-5}$, a batch size of 256, and a temperature parameter $\tau=1.0$.
@@
After training the representations, we freeze the encoders and, for each bias setting, train both a linear classifier (logistic regression) and a nonlinear classifier (a two-layer MLP with ReLU activation) for each discrete latent factor. Similarly, for continuous latent factors, we train both a linear regressor and a nonlinear regressor (a two-layer MLP with ReLU activation). All classifiers and regressors are trained for 10,00010{,}000 steps using the evaluation subset.
@@ F.2 Additional Results
We now turn to the analysis of the text representations learned by MMCL. As shown in Figure˜19, we observe patterns similar to those found in the image modality. In particular, discrete latent semantic variables—such as shape, x_pos, and y_pos—are reliably identified, provided they are unbiased and consistently aligned across modalities. A notable case is s_pos, which is a continuous latent factor in both modalities but is mapped to only five discrete tokens in the text observations, rendering the text generation process non-invertible. Despite this lossy transformation, the model achieves a relatively high $R^{2}$, suggesting that the learned text representations remain strongly influenced by the alignment objective, even when semantic information is partially lost.
@@
We assess the linear separability of the learned image and text representations by evaluating the performance of linear classifiers and regressors trained to predict each latent factor. As shown in Figure˜20 and Figure˜21, and in comparison to the nonlinear prediction results, the findings indicate that not all latent semantic variables are linearly embedded in the representations learned through MMCL. In particular, the semantic factor color consistently demonstrates poor linear predictability, suggesting that it is encoded nonlinearly in both modalities. Conversely, factors such as x_pos and y_pos exhibit strong linear separability in certain settings, indicating that these semantic variables are more directly captured in the latent space of both the image and text encoders. Overall, these results suggest that the linearity of the learned representations is factor-dependent and shaped by both modality-specific encoding strategies and the underlying structure of the input data—highlighting an important direction for future investigation.
@@
Appendix G Details on OpenCLIP Case Study
G.1 Concepts Taxonomy and Data Collection
Concepts Taxonomy.
To validate the representations of pretrained OpenCLIP models, we curated a set of 146 concepts organized into 15 distinct concept groups. Each concept group is treated as a separate solution space, and we aim to ensure that the concepts within each group are mutually exclusive. Table˜12 presents all the concepts used in the case study. For simplicity, abbreviations are used in the main text, with their full meanings provided in the concept group column.
Table 12: Taxonomy of concept categories with abbreviations, grouped by conceptual relevance.
Group Explanation Concept Name
Animal animals dog, cat, horse, bird, elephant, giraffe, cow, zebra, rabbit, duck
Clothing clothing shirt, pants, dress, shoes, hat, jacket, skirt, tie, hoodie, socks
Color colors red, blue, green, yellow, black, white, orange, pink, gray, brown
Food food pizza, burger, sandwich, salad, cake, coffee, tea, beer, ice cream, noodles
Object object chair, table, phone, laptop, car, bottle, bag, cup, backpack, television
Role roles chef, teacher, athlete, doctor, engineer, artist, pilot, firefighter, police officer, lawyer
Scene scenes beach, kitchen, forest, street, park, office, bedroom, classroom, stadium, playground
Vehicle vehicles bus, truck, airplane, train, motorcycle, bicycle, boat, van, taxi, scooter
Weather weathers sandstorm, drought, rain, snow, fog, rainbow, storm, moonlight, overcast, frost
Texture texture glossy, matte, rough, smooth, fuzzy, silky, grainy, wrinkled, slippery, furry
POV viewpoint close-up, wide shot, top-down, low angle, high angle, aerial, shallow-depth, long exposure
Emot. mental state tired, focused, surprised, proud, shy, bored, confused, excited, thoughtful, nervous
Postproc. post-procssing hdr, desaturated, color-filtered, over-sharpened, saturated, sepia, heavily edited, bokeh, tilt-shift
Trait moral bias easy-going, cold-hearted, evil, criminal, immoral, pure, corrupt, hero, cheater, deserving
Stere. stereotypes exotic look, ethnic look, slacker, aggressive look, gangster look, terrorist look, illegal look, privileged, ghetto look
Data Collection.
We collected up to 200 images per concept by using the concept names as search queries on Flickr, filtering for content licensed under Creative Commons. To ensure relevance, all images were manually reviewed, and those deemed unrelated to the queried concept were removed. The remaining images were retained for subsequent experiments. Table˜13 summarizes the number of images retained for each concept. It is important to note that some concept groups, such as those related to stereotypes, include terms that may reflect societal biases. We do not endorse or perpetuate these biases; rather, these concepts are included solely to assess whether pretrained models encode or reproduce such biased associations in their representations.
Table 13: Concept-wise image counts grouped by concept category for zero-shot evaluation.
Group Concept Count Group Concept Count Group Concept Count
Animal dog 124 Animal cat 112 Animal horse 109
Animal bird 106 Animal elephant 98 Animal giraffe 90
Animal cow 88 Animal zebra 81 Animal rabbit 77
Animal duck 65 Clothing shirt 113 Clothing pants 104
Clothing dress 101 Clothing shoes 96 Clothing hat 92
Clothing jacket 87 Clothing skirt 84 Clothing tie 78
Clothing hoodie 72 Clothing socks 68 Color red 130
Color blue 127 Color green 119 Color yellow 114
Color black 111 Color white 109 Color orange 97
Color pink 91 Color gray 88 Color brown 82
Food pizza 118 Food burger 113 Food sandwich 107
Food salad 101 Food cake 95 Food coffee 93
Food tea 90 Food beer 85 Food ice cream 82
Food noodles 78 Emot. tired 58 Emot. focused 55
Emot. surprised 51 Emot. proud 49 Emot. shy 45
Emot. bored 44 Emot. confused 42 Emot. excited 40
Emot. thoughtful 39 Emot. nervous 36 Trait easy-going 29
Trait cold-hearted 27 Trait evil 25 Trait criminal 24
Trait immoral 23 Trait pure 22 Trait corrupt 20
Trait hero 18 Trait cheater 17 Trait deserving 15
Object chair 120 Object table 115 Object phone 112
Object laptop 110 Object car 108 Object bottle 105
Object bag 100 Object cup 97 Object backpack 93
Object television 90 Postproc. hdr 33 Postproc. desaturated 32
Postproc. color-filtered 31 Postproc. over-sharpened 30 Postproc. saturated 28
Postproc. sepia 27 Postproc. heavily edited 26 Postproc. bokeh 24
Postproc. tilt-shift 22 Role chef 88 Role teacher 86
Role athlete 85 Role doctor 83 Role engineer 80
Role artist 78 Role pilot 74 Role firefighter 71
Role police officer 68 Role lawyer 66 Scene beach 120
Scene kitchen 115 Scene forest 110 Scene street 107
Scene park 104 Scene office 102 Scene bedroom 98
Scene classroom 95 Scene stadium 92 Scene playground 89
Stere. exotic look 21 Stere. ethnic look 20 Stere. slacker 19
Stere. aggressive look 18 Stere. gangster look 17 Stere. terrorist look 16
Stere. illegal look 15 Stere. privileged 14 Stere. ghetto look 13
Texture glossy 47 Texture matte 45 Texture rough 43
Texture smooth 41 Texture fuzzy 40 Texture silky 38
Texture grainy 36 Texture wrinkled 34 Texture slippery 33
Texture furry 30 Vehicle bus 106 Vehicle truck 104
Vehicle airplane 102 Vehicle train 100 Vehicle motorcycle 97
Vehicle bicycle 93 Vehicle boat 89 Vehicle van 86
Vehicle taxi 83 Vehicle scooter 80 POV close-up 55
POV wide shot 52 POV top-down 50 POV low angle 48
POV high angle 47 POV aerial 44 POV shallow-depth 42
POV long exposure 39 Weather sandstorm 66 Weather drought 63
Weather rain 61 Weather snow 58 Weather fog 56
Weather rainbow 53 Weather storm 50 Weather moonlight 47
Weather overcast 45 Weather frost 42
G.2 Details on Caption Frequency Statistics
To assess caption coverage frequency for each concept, we first enumerated plausible synonyms for every concept name. We then computed the percentage of captions in the LAION-400M dataset that included either the original concept term or one of its synonyms. Figure˜9 in §˜5.4 presents the group-wise average coverage rates.
To explore intra-group variability, Figure˜22 provides a more granular view of concept-level coverage. It reveals that even within a single concept group, the degree of caption coverage can vary substantially across individual concepts. As expected, commonly observable concepts, such as object types, colors, and food items, tend to appear more frequently in captions. In contrast, abstract or subjective concepts, such as mental states, post-processing artifacts, moral judgments, and social stereotypes, are significantly underrepresented. This discrepancy may arise from several factors: ethical concerns leading annotators to avoid certain terms, unconscious omission of nuanced details, or general challenges in describing abstract visual content.
Figure 22: Statistics of concept-level coverage within each group in LAION-400M captions.
G.3 Zero-Shot Evaluation Details
To probe the representations of pretrained OpenCLIP models, we evaluate two variants trained on the LAION-400M dataset: ViT-B-32 and ViT-L-14. Using the collected image samples, we perform zero-shot analysis by employing the concept names as text prompts. Each concept group is treated as a distinct solution space. To mitigate the impact of sampling bias in the test dataset, we report macro-averaged F1 scores at the group level.
As shown in Figure˜10, concept groups that are frequently captioned—defined here as those with coverage rates exceeding 0.15%, are generally well represented in the multimodal contrastive learning (MMCL) models, with F1 scores above 65%. In contrast, concept groups that are severely under-captioned, often due to selection bias or ethical omission, exhibit poor representation, with F1 scores falling below 40%. The results are consistent across different model scales, further reinforcing our theoretical findings, particularly the impact of selection bias on the learned representations.
Figure˜23 presents the intra-group confusion matrix from the zero-shot evaluation using the OpenCLIP ViT-B-32 model trained on the LAION-400M dataset. As clearly shown, well-captioned concept groups exhibit strong diagonal patterns, indicating accurate predictions—whereas concept groups affected by selection bias suffer from substantial misclassification.
This misclassification has multifaceted implications. For certain under-captioned yet potentially valuable concepts, such as Texture (texture) and Emot. (mental state), the poor performance suggests a need for more comprehensive and semantically rich captioning pipelines in multimodal contrastive learning (MMCL) pretraining. Conversely, for sensitive concepts like Trait (moral judgement) and Stere. (stereotypes), which ideally should not be encoded, this under-representation may be desirable. In these cases, the inherent biases in the captioning process may function as epistemic filters that implicitly align model representations with human ethical standards.
Figure 23: Confusion matrix for zero-shot prediction using OpenCLIP ViT-B-32 of each group.
Appendix H Computation Resources
Appendix I Discussions